(2 points) Listed below are the lead concentrations in ug/g measured in different traditional medicines. Use a 0.05 significance level to test the claim that the mean lead concentration for all such medicin es is less than 14 μg/g. 2.5 14.5 17.5 7.514.5 18.5 10.5 20 a. Determine the test statistic (Blank 1). Round to two decimal places.) b. Determine the P-value (Blank 2). Round to three decimal places.) Blank # 1 Blank # 2 Question 26 (3 points) (Answer Questions 24 through 26.) (3 points) Listed below are the lead concentrations in ug/g measured in different traditional medicines. Use a 0.05 significance level to test the claim that the mean lead concentration for all such medicines is less than 14 μg/g. 2.5 14 17.5 7.5 14 18 10.5 20 State the final conclusion that addresses the original claim. Choose the correct answer for each enumerated blank space in the two incomplete sentences below. You can either copy and paste the two sentences and the correct answer (CAPITAL LETTERS) for each enumerated blank space, or you can copy and paste the number with the respective correct answer in the response box. In case you cannot copy and paste, you just need to type the correct answers with their respective number.) 1. REJECT, FAIL TO REJECT 2. SUFFICIENT, NOT SUFFICIENT 3. LESS THAN, GREATER THAN, EQUAL TO Ho. There is (2) mean lead concentration for all such medicines is (3) evidence to conclude that the 14 Hg/g.
Answer #1

First enter Data into EXCEL

We have to find the sample mean.

Excel command is =AVERAGE(Select data)

sample mean = 11.35

Now we have to find sample standard deviation.

Excel command is =STDEV(Select data)

standard deviation = s = 6.562

n = 10

11.35

S =6.562

claim : The mean lead concentration for such medicines is less than 14 ug/g

Null and alternative hypothesis is

H0 : u = 14

H1 : u < 14

Level of significance = 0.05

Here population standard deviation is not known so we use t-test statistic.

Test statistic is

Vn

11.35 - 14 10

-2.65 6.562 10

一2.65 2.075 t=

1.277

t _-1.28

Degrees of freedom = n - 1 =10-1=9

a-0.05

P-value = 0.1168   ( using t table)

P-value =0.117

Blank # 1 = t _-1.28

Blank # 2 = P-value =0.117

Quetion 26

P-value >alpha ,Failed to Reject Ho

1.FAIL TO REJECT

2. NOT SUFFICIENT

3.LESS THAN

(1) FAIL TO REJECT Ho.There is (2) NOT SUFFICIENT evidence to conclude that the mean lead concentration for such medicines is LESS THAN14 ug/g